# Two-phase flows through porous media described by a Cahn–Hilliard–Brinkman model with dynamic boundary conditions

### v1

@@ 2.2   General assumptions
    The viscosity function $\nu:{\mathbb{R}}\to{\mathbb{R}}$ and the friction parameter $\gamma:{\mathbb{R}}\to{\mathbb{R}}$ are continuous, bounded and uniformly positive. Namely, there exist positive constants $\nu_{1}$, $\nu_{2}$, $\gamma_{1}$ and $\gamma_{2}$ such that
    $0<\nu_{1}\leq\nu(r)\leq\nu_{2},\quad 0<\gamma_{1}\leq\gamma(r)\leq\gamma_{2}\quad\text{for all $r\in{\mathbb{R}}.$}$
    Furthermore, the permeability function $\lambda:{\mathbb{R}}\to{\mathbb{R}}$ is continuous, bounded and nonnegative, that is, there exists a nonnegative constant $\lambda_{1}$ such that
    $0\leq\lambda(r)\leq\lambda_{1}\quad\text{for all $r\in{\mathbb{R}}.$}$
@@ 2.3   Preliminaries
Throughout the paper, we will frequently use the following bulk-surface Poincaré inequality:
@@
There exists a constant $C_{P}^{\prime}$ depending only on $\Omega$ such that
$\|u\|_{H}\leq C_{P}^{\prime}\left(\|\nabla u\|_{{\bf H}}+\|u\|_{{H}_{\Gamma}}\right)\quad\text{for all $u\in V$.}$ (2.3)
A proof of this inequality can be found, for instance, in [61, Chapter II, Section 1.4].
@@
Inequality (2.4) follows from an interpolation result shown in [62, Sec 4.3.1, Theorem 1 and Remark 1], whereas (2.5) follows from an interpolation result presented in [63, Sec 7.4.5, Remark 2].
@@ 2.4   The Cahn–Hilliard–Brinkman system with regular potentials
First, we present our mathematical results for system (1) in the case of regular double-well potentials $F$ and $G$. As mentioned above, we simply set $\epsilon=\epsilon_{\Gamma}=1$, as the exact values of these interface parameters do not have any impact on the mathematical analysis as long as they are positive.
@@ 2.4.2   Definition of weak solutions for regular potentials
Let $K\geq 0$ be arbitrary. Suppose that (A1)(A3) and (R1) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be any initial data. A quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ is called a weak solution of the Cahn–Hilliard–Brinkman system (1) if the following conditions are fulfilled:
@@
[…]
@@
[…]
@@
[…]
@@ 2.4.3   Existence of a weak solution in the case $K>0$
We first show the existence of a weak solution to the Cahn–Hilliard–Brinkman system (1) in the case $K>0$.
@@
Let $K>0$ be arbitrary. Suppose that (A1)(A3) and (R1) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be any initial data. Then, the Cahn–Hilliard–Brinkman system (1) possesses at least one weak solution $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ in the sense of Definition 2.3, which further satisfies $(\mu,\theta)\in L^{4}(0,T;{\cal H})$.
Let us now assume that the domain $\Omega$ is of class $C^{\ell}$ with $\ell\in\{2,3\}$. If $d=3$, we further assume $p<6$, and if $\ell=3$, we further assume that (R2) holds. Then, we have the additional regularities
@@ 2.4.4   The limit $K\to 0$ and existence of a weak solution in the case $K=0$
We now investigate the limit $K\to 0$ in which the boundary condition (1.1g) formally tends to the Dirichlet condition $\psi=\varphi|_{\Gamma}$ almost everywhere on $\Sigma$. In the following theorem, we send $K\to 0$ in system (1) to prove the existence of a weak solution to (1) in the case $K=0$, and we further specify the convergence properties of this asymptotic limit.
@@
${\boldsymbol{v}}^{K_{n}}$ $\to{\boldsymbol{v}}^{0}$ weakly in $\quad\text{weakly in $L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})$},$ (2.17a)
${\boldsymbol{v}}^{K_{n}}|_{\Gamma}$ $\to{\boldsymbol{v}}^{0}|_{\Gamma}$ weakly in $\quad\text{weakly in $L^{2}(0,T;\mathbf{H}_{\Gamma})$},$ (2.17b)
[…]
strongly in C0([0,T];Hs(Ω)), and a.e. in $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~{}$Q$}},$ (2.17c)
[…]
strongly in C0([0,T];Hs(Γ)), and a.e. on $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~{}$\Sigma$}},$ (2.17d)
$\mu^{K_{n}}$ $\to\mu^{0}$ weakly in $\quad\text{weakly in $L^{2}(0,T;V)$},$ (2.17e)
$\theta^{K_{n}}$ $\to\theta^{0}$ weakly in $\quad\text{weakly in $L^{2}(0,T;{V}_{\Gamma})$},$ (2.17f)
$\varphi^{K_{n}}|_{\Gamma}-\psi^{K_{n}}$ $\to 0$ strongly in L(0,T;HΓ), and a.e. on $\quad\text{strongly in $L^{\infty}(0,T;{H}_{\Gamma})$, and {a.e.\ on~{}$\Sigma$}},$ (2.17g)
as $n\to\infty$ along a non-relabeled subsequence. Moreover, the limit $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ is a weak solution of the Cahn–Hilliard–Brinkman model (1) in the sense of Definition 2.3 with $K=0$.
@@
Remark 2.6.
The additional assumption $p\leq 4$ in the second paragraph of the above theorem was merely imposed to avoid unnecessary technicalities in the proof for additional regularity of the phase-fields. However, it should be possible to obtain the same regularities also for $p\in(4,6)$. This could be established by approximating the potential $F$ by a sequence $(F_{k})_{k\in{\mathbb{N}}}$ of potentials satisfying (R1) with $p\leq 4$. Then, proceeding similarly as in the proof of Theorem 2.4 (see Subsection 3.1.5) additional uniform bounds on the corresponding weak solutions $({\boldsymbol{v}}_{k},\varphi_{k},\mu_{k},\psi_{k},\theta_{k})$ would have to be derived to obtain the weak convergences $(\varphi_{k},\psi_{k})\to(\varphi,\psi)$ in $L^{2}\big{(}0,T;H^{2}(\Omega)\times H^{2}(\Gamma)\big{)}$ in case $\ell\in\{2,3\}$, and $(\varphi_{k},\psi_{k})\to(\varphi,\psi)$ in $L^{2}\big{(}0,T;H^{3}(\Omega)\times H^{3}(\Gamma)\big{)}$ in case $\ell=3$, which imply the desired regularities.
@@ 2.4.5   Stability and uniqueness of the weak solution in the general case $K\geq 0$
    $p\in\begin{cases}[3,\infty)&\text{if $d=2$,}\\ [3,6)&\text{if $d=3$ and $K>0$,}\\ [3,4]&\text{if $d=3$ and $K=0$,}\end{cases}\qquad\text{and}\qquad q\in[3,\infty)$ (2.21)
@@
We point out that (Renumi) implies (R1) and (R2) with $p$ and $q$ being chosen as in (2.21). The restriction $p\in[3,4]$ in the case $d=3$ and $K=0$ is imposed since in Theorem 2.5, higher regularity properties for the phase-fields were established only for $p\leq 4$. However, as pointed out in Remark 2.6, it should be possible to relieve this restriction such that also in the case $d=3$ and $K=0$, merely $p\in[3,6)$ would have to be imposed in (Renumi).
@@
Suppose that (A1) and (Renumi) are fulfilled and let $K\geq 0$ be arbitrary. In addition to (A2) and (A3), we further assume that $\gamma$ and $\lambda$ are Lipschitz continuous functions and that the functions $\nu$, $M_{\Omega}$ and $M_{\Gamma}$ reduce to positive constants denoted by the same symbols. For any $i\in\{1,2\}$, let $(\varphi_{0,i},\psi_{0,i})\in{\cal V}_{K}$ be any pair of initial data, and let $({\boldsymbol{v}}_{i},\varphi_{i},\mu_{i},\psi_{i},\theta_{i})$ be a corresponding weak solution in the sense of Definition 2.3. Then, the stability estimate
@@ 2.5   The Cahn–Hilliard–Brinkman system with singular potentials
We now consider the system (1) for a general class of singular potentials. For those, we manage to establish just existence of weak solutions due to the lower regularity at disposal. Recall that $\epsilon=\epsilon_{\Gamma}=1$ as mentioned above.
@@ 2.5.1   Assumptions for singular potentials
    $\beta:=\partial\widehat{\beta}\quad\text{and}\quad\beta_{\Gamma}:=\partial{\widehat{\beta}_{\Gamma}}.$
    Moreover, we suppose that $\widehat{\pi},{\widehat{\pi}}_{\Gamma}\in C^{1}({\mathbb{R}})$ with Lipschitz continuous derivatives $\pi:=\widehat{\pi}^{\prime}$ and $\pi_{\Gamma}:={\widehat{\pi}}_{\Gamma}^{\prime}$.
    We point out that $\beta$ and ${\beta_{\Gamma}}$ are maximal monotone graphs in ${\mathbb{R}}\times{\mathbb{R}}$ whose effective domains are denoted by $D(\beta)$ and $D(\beta_{\Gamma})$, respectively. In particular, as 00 is a minimum point of both $\widehat{\beta}$ and ${\widehat{\beta}}_{\Gamma}$, it turns out that $0\in\beta(0)$ and $0\in{\beta_{\Gamma}}(0)$. Finally, we denote by $\beta^{\circ}$ the minimal section of the graph $\beta$, which is defined as

### v2

@@ 2.2  General assumptions
    The viscosity function $\nu:{\mathbb{R}}\to{\mathbb{R}}$, the permeability function $\lambda:{\mathbb{R}}\to{\mathbb{R}}$, and the friction parameter $\gamma:{\mathbb{R}}\to{\mathbb{R}}$ are continuous and nonnegative. There further exist constants $\nu_{1},\nu_{2},\lambda_{2},\gamma_{2}>0$ as well as $\gamma_{1}\geq 0$ such that
    $0<\nu_{1}\leq\nu(r)\leq\nu_{2},\quad 0\leq\lambda(r)\leq\lambda_{2},\quad 0\leq\gamma_{1}\leq\gamma(r)\leq\gamma_{2}\quad\text{for all $r\in{\mathbb{R}}.$}$
    Moreover, we assume that one of the following conditions holds:
    1. (A3.1)
      It holds $\gamma_{1}>0$.
    2. (A3.2)
      The domain $\Omega$ has the following property:
[…]
@@ 2.3  Preliminaries
In our mathematical analysis, we will need the following versions of Korn’s inequality.
@@
Suppose that (A1) holds.
  1. (a)
    There exists a constant $C_{\mathrm{Korn}}$ depending only on $\Omega$ such that
    $\|\nabla{\boldsymbol{v}}\|_{{\mathbb{H}}}\leq C_{\mathrm{Korn}}\left(\|\mathbf{D}{\boldsymbol{v}}\|_{{\mathbb{H}}}+\|{\boldsymbol{v}}\|_{\mathbf{H}_{\Gamma}}\right)\quad\text{for all ${\boldsymbol{v}}\in{\bf{V}}$.}$ (2.4)
  2. (b)
    If $\Omega$ additionally fulfills the condition (2.3), then there exists a constant $C_{\mathrm{Korn}}^{*}$ depending only on $\Omega$ such that
    $\|\nabla{\boldsymbol{v}}\|_{{\mathbb{H}}}\leq C_{\mathrm{Korn}}^{*}\|\mathbf{D}{\boldsymbol{v}}\|_{{\mathbb{H}}}\quad\text{for all ${\boldsymbol{v}}\in{\bf{V}}$.}$ (2.5)
Both (a) and (b) can be found in [3, Appendix]. In the three-dimensional case, (b) had already been established before in [69, Theorem 3.5]. As illustrated in [3], the two-dimensional version can be proved analogously.
Moreover, besides the standard Poincaré–Wirtinger inequality in $\Omega$, we need a Poincaré type inequality on $\Gamma$.
Lemma 2.2.
Suppose that (A1) holds. Then, there exists a constant $C_{\mathrm{P}}\geq 0$ depending only on $\Omega$ such that
$\|v-{\langle v\rangle}_{\Gamma}\|_{{H}_{\Gamma}}\leq C_{P}\|\nabla v\|_{\mathbf{H}_{\Gamma}}\quad\text{for all $v\in{V}_{\Gamma}$.}$ (2.6)
This result follows directly from the bulk-surface Poincaré inequality established in [61, Lemma A.1] (with the function in the bulk being chosen as $u\equiv 0$ and the parameters being chosen as $K=2$, $\alpha=\beta=1$).
@@
Inequality (2.7) follows from an interpolation result shown in [77, Sec 4.3.1, Theorem 1 and Remark 1], whereas (2.8) follows from an interpolation result presented in [78, Sec 7.4.5, Remark 2].
@@ 2.4  The Cahn–Hilliard–Brinkman system with regular potentials
First, we present our mathematical results for system (1.1) in the case of regular double-well potentials $F$ and $G$. As mentioned above, we simply set $\epsilon=\epsilon_{\Gamma}=1$, as the exact values of these interface parameters do not have any impact on the mathematical analysis as long as they are positive.
@@ 2.4.2 Definition of weak solutions for regular potentials
Let $K\geq 0$ be arbitrary. Suppose that (A1)(A3) and (R1) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be any initial data. A quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ is called a weak solution of the Cahn–Hilliard–Brinkman system (1.1) if the following conditions are fulfilled:
@@
    $\begin{aligned} &2\int_{\Omega}\nu(\varphi)\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{w}}+\int_{\Omega}\lambda(\varphi){\boldsymbol{v}}\cdot{\boldsymbol{w}}+\int_{\Gamma}\gamma(\psi){\boldsymbol{v}}\cdot{\boldsymbol{w}}\\ &\quad=-\int_{\Omega}\varphi\nabla\mu\cdot{\boldsymbol{w}}-\int_{\Gamma}\psi\nabla_{\Gamma}\theta\cdot{\boldsymbol{w}},\end{aligned}$ (2.16a)
@@
    $\begin{aligned} \int_{\Omega}\mu\eta+\int_{\Gamma}\theta{\eta_{\Gamma}}&=\int_{\Omega}\nabla\varphi\cdot\nabla\eta+\int_{\Omega}F^{\prime}(\varphi)\eta+\int_{\Gamma}\nabla_{\Gamma}\psi\cdot\nabla_{\Gamma}{\eta_{\Gamma}}\\ &\quad+\int_{\Gamma}G^{\prime}(\psi){\eta_{\Gamma}}+\sigma(K)\int_{\Gamma}(\psi-\varphi)({\eta_{\Gamma}}-\eta)\end{aligned}$ (2.16d)
@@
    $\varphi(0)=\varphi_{0}\quad\hbox{a.e.\ in~$\Omega$},\quad\psi(0)=\psi_{0}\quad\hbox{a.e.\ on~$\Gamma$}.$
@@
Remark 2.5.
We point out that the pressure $p$ does not appear in the weak formulation (2.16a) as the test functions are chosen to be divergence-free. However, provided that $\Omega$ is connected, the pressure can be reconstructed in the following way. Suppose that $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ is a weak solution in the sense of Definition 2.4. We define
$\mathcal{F}:W_{\bf n}\to{\mathbb{R}},\quad\mathcal{F}(\Phi)$ $=\int_{\Omega}2\nu(\varphi)\mathbf{D}{\boldsymbol{v}}:\mathbf{D}\nabla\Phi+\int_{\Omega}\lambda(\varphi){\boldsymbol{v}}\cdot\nabla\Phi+\int_{\Gamma}\gamma(\psi){\boldsymbol{v}}\cdot\nabla\Phi$
$\quad+\int_{\Omega}\varphi\nabla\mu\cdot\nabla\Phi+\int_{\Gamma}\psi\nabla_{\Gamma}\theta\cdot\nabla\Phi\,.$
Note that in the integrals over $\Gamma$, we can actually replace $\nabla\Phi$ by $\nabla_{\Gamma}\Phi$ as $\boldsymbol{v}$} and $\nabla_{\Gamma}\theta$ are tangential vector fields, i.e., their normal component is zero. In view of the regularities in Definition 2.4(i), it is straightforward to check that $\mathcal{F}\in W_{\bf n}^{*}$. Hence, according to [1, Lemma 3.6.1], there exists a unique function $p\in L^{2}(\Omega)$ such that
$\int_{\Omega}p\,\Delta\Phi=\mathcal{F}(\Phi)\quad\text{for all $\Phi\in W_{\bf n}$.}$ (2.18)
Let now $\widehat{\boldsymbol{w}}\in{\bf{V}}_{\bf n}$ be arbitrary. As $\Omega$ is additionally assumed to be connected, there exists a Leray decomposition $\widehat{\boldsymbol{w}}={\boldsymbol{w}}+\nabla\Phi$} with ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$ and $\Phi\in W_{\bf n}$ (see, e.g., [15, Theorem IV.3.5.]). In particular, we thus have $\mathop{\rm div}\nolimits\widehat{\boldsymbol{w}}=\mathop{\rm div}\nolimits(\nabla\Phi)=\Delta\Phi$. Hence, combining (2.16a) and (2.18), we conclude that the variational formulation
$2\int_{\Omega}\nu(\varphi)\mathbf{D}{\boldsymbol{v}}:\mathbf{D}\widehat{\boldsymbol{w}}-\int_{\Omega}p\mathop{\rm div}\nolimits\widehat{\boldsymbol{w}}+\int_{\Omega}\lambda(\varphi){\boldsymbol{v}}\cdot\widehat{\boldsymbol{w}}+\int_{\Gamma}\gamma(\psi){\boldsymbol{v}}\cdot\widehat{\boldsymbol{w}}$
$\quad=-\int_{\Omega}\varphi\nabla\mu\cdot\widehat{\boldsymbol{w}}-\int_{\Gamma}\psi\nabla_{\Gamma}\theta\cdot\widehat{\boldsymbol{w}}$
holds for all $\widehat{\boldsymbol{w}}\in{\bf{V}}_{\bf n}$. We have thus reconstructed the pressure $p\in L^{2}(\Omega)$.
@@ 2.4.3 Existence of a weak solution in the case $K>0$
We first show the existence of a weak solution to the Cahn–Hilliard–Brinkman system (1.1) in the case $K>0$.
@@
Let $K>0$ be arbitrary. Suppose that (A1)(A3) and (R1) are fulfilled and let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ be any initial data. Then, the Cahn–Hilliard–Brinkman system (1.1) possesses at least one weak solution $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ in the sense of Definition 2.4, which further satisfies $(\mu,\theta)\in L^{4}(0,T;{\cal H})$.
@@ 2.4.4 The limit $K\to 0$ and existence of a weak solution in the case $K=0$
We now investigate the limit $K\to 0$ in which the boundary condition (1.1g) formally tends to the Dirichlet condition $\psi=\varphi|_{\Gamma}$ almost everywhere on $\Sigma$. In the following theorem, we send $K\to 0$ in system (1.1) to prove the existence of a weak solution to (1.1) in the case $K=0$, and we further specify the convergence properties of this asymptotic limit.
@@
${\boldsymbol{v}}^{K_{n}}$ $\to{\boldsymbol{v}}^{0}$ $\quad\text{weakly in $L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})$},$ (2.21a)
${\boldsymbol{v}}^{K_{n}}|_{\Gamma}$ $\to{\boldsymbol{v}}^{0}|_{\Gamma}$ $\quad\text{weakly in $L^{2}(0,T;\mathbf{H}_{\Gamma})$},$ (2.21b)
$\varphi^{K_{n}}$ $\to\varphi^{0}$ $\quad\text{weakly-${}^{*}$ in $L^{\infty}(0,T;V)$, weakly in $H^{1}(0,T;V^{*})$},$
$\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~$Q$}},$ (2.21c)
$\psi^{K_{n}}$ $\to\psi^{0}$ $\quad\text{weakly-${}^{*}$ in $L^{\infty}(0,T;{V}_{\Gamma})$, weakly in $H^{1}(0,T;{V}_{\Gamma}^{*})$},$
$\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~$\Sigma$}},$ (2.21d)
$\mu^{K_{n}}$ $\to\mu^{0}$ $\quad\text{weakly in $L^{2}(0,T;V)$},$ (2.21e)
$\theta^{K_{n}}$ $\to\theta^{0}$ $\quad\text{weakly in $L^{2}(0,T;{V}_{\Gamma})$},$ (2.21f)
$\varphi^{K_{n}}|_{\Gamma}-\psi^{K_{n}}$ $\to 0$ $\quad\text{strongly in $L^{\infty}(0,T;{H}_{\Gamma})$, and {a.e.\ on~$\Sigma$}},$ (2.21g)
as $n\to\infty$ along a non-relabeled subsequence. Moreover, the limit $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ is a weak solution of the Cahn–Hilliard–Brinkman model (1.1) in the sense of Definition 2.4 with $K=0$.
@@ 2.4.5 Stability and uniqueness of the weak solution in the general case $K\geq 0$
    $p\in\begin{cases}[3,\infty)&\text{if $d=2$,}\\ [3,4]&\text{if $d=3$,}\end{cases}\qquad\text{and}\qquad q\in[3,\infty),$ (2.25)
@@
We point out that (Renumi) implies (R1) and (R2) with $p$ and $q$ being chosen as in (2.25).
@@
Suppose that (A1) and (Renumi) are fulfilled with $\Omega$ being of class $C^{3}$, and let $K\geq 0$ be arbitrary. In addition to (A2) and (A3), we further assume that $\gamma$ and $\lambda$ are Lipschitz continuous and that the functions $\nu$, $M_{\Omega}$ and $M_{\Gamma}$ reduce to positive constants denoted by the same symbols. For any $i\in\{1,2\}$, let $(\varphi_{0,i},\psi_{0,i})\in{\cal V}_{K}$ be any pair of initial data, and let $({\boldsymbol{v}}_{i},\varphi_{i},\mu_{i},\psi_{i},\theta_{i})$ be a corresponding weak solution in the sense of Definition 2.4. Then, the stability estimate
@@ 2.5  The Cahn–Hilliard–Brinkman system with singular potentials
We now consider the system (1.1) for a general class of singular potentials. For those, we manage to establish just existence of weak solutions due to the lower regularity at disposal. Recall that $\epsilon=\epsilon_{\Gamma}=1$ as mentioned above.
@@ 2.5.1 Assumptions for singular potentials
    $\beta:=\partial\widehat{\beta}\quad\text{and}\quad\beta_{\Gamma}:=\partial{\widehat{\beta}_{\Gamma}},$
    where $\partial$ indicates the subdifferential of the respective function. Moreover, we suppose that $\widehat{\pi},{\widehat{\pi}}_{\Gamma}\in C^{1}({\mathbb{R}})$ with Lipschitz continuous derivatives $\pi:=\widehat{\pi}^{\prime}$ and $\pi_{\Gamma}:={\widehat{\pi}}_{\Gamma}^{\prime}$.
    We point out that $\beta$ and $\beta_{\Gamma}$} are maximal monotone graphs in $\mathbb{R}$$\times{\mathbb{R}}$ whose effective domains are denoted by $D(\beta)$ and $D(\beta_{\Gamma})$, respectively. In particular, as 0 is a minimum point of both $\widehat{\beta}$ and ${\widehat{\beta}}_{\Gamma}$, it turns out that $0\in\beta(0)$ and $0\in{\beta_{\Gamma}}(0)$. Finally, we denote by $\beta^{\circ}$ the minimal section of the graph $\beta$, which is defined as